OpenStudy (elonasushchik):

Hey!!!! Consider the function f(x) = 4x2 + 5. Find the value of f(x) when x = 2 and select the correct answer below. 13 21 69 89

OpenStudy (steve816):

Hey, to find the value when x=2, simply plug in 2 to your equation!

OpenStudy (elonasushchik):

okay then..

OpenStudy (steve816):

\[f(x)=4(2)^2+5\]

OpenStudy (steve816):

\[f(x)=4(4)+5\] \[f(x)=16+5\] \[f(x)=21\]

OpenStudy (steve816):

So your answer would be the 2nd choice. Hope this helps!

OpenStudy (elonasushchik):

I knew it was B. Thank you Sooo much steve for justifying my answer and taking your time!!! @steve816

OpenStudy (steve816):

No problem!

OpenStudy (elonasushchik):

:)